Mdm Yip had some oranges. She packed 7/9 of the oranges into 42 bags and the rest into 4 boxes. Each box had 34 oranges than each bag.

A) What fraction of the oranges were packed into 4 boxes?

B) How many oranges did Mdm Yip pack in each bag?


1
Expert's answer
2021-03-18T13:37:59-0400

A) The fraction of the oranges that were packed into 4 boxes is:

1 - 7/9 = 2/9.

B) If we suppose that there is a mistake in the text and each box had 34 oranges MORE than each bag, and x is the number of oranges that Mdm Yip packed in each bag, then:

"42x\u00f77\/9 = 4(x + 34)\u00f72\/9,"

"42x = (7\/9\u00f72\/9)\u00d74(x + 34),"

42x =14x + 476,

28x = 476,

x = 17 oranges.